OpenStudy (anonymous):

solve the system by method of substitution: -x+y=-9 x^2+y^2=81

OpenStudy (anonymous):

solve the first equation for y -x + y = -9 y = x+9 plug x+9 in for y in the next equation solve the quadratic formula for x.

OpenStudy (anonymous):

k thanks!!!

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

i now have x^2+18x+81=81.will the 81's cancel each other out? @tomo

OpenStudy (anonymous):

x^2+y^2=81 x^2 + (x+9)^2 = 81 x^2 + x^2 + 18x + 81 = 81 subtract 81 from both sides 2x^2 + 18x = 0 2x^2 = -18x x^2 = -9x x^2 + 9x = 0 x*(x+9) = 0 // one solution x = 0 // other solution x+9 = 0 x = -9